# Primary gravitational waves at high frequencies I: Origin of suppression in the power spectrum

### v1

@@ 3.2 PS during matter domination
Since $a_{0}$ and $H_{0}$ are the scale factor and the Hubble parameter today, $k_{0}/a_{0}=H_{0}$ is the largest wave number of observational interest. On substituting the above expressions for $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ in the PS (3.16) and considering the limit $z\gg 1$, which corresponds to $k_{0}\ll k\ll k_{\mathrm{eq}}$, we obtain that
@@
Figure 2: The PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ evaluated today, i.e. at $\eta_{0}$, has been plotted for the case of instantaneous transition from de Sitter inflation to the epoch of radiation domination and subsequently to the epoch of matter domination [cf. Eq. (3.2)]. We have plotted the PS for the three values of $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}$ we had plotted in the previous figure. Note that, in addition to the different behavior of the PS in the three regimes mentioned in the previous figure, there arises a fourth regime when $x_{\mathrm{eq}}\ll 1$ and $z\gg 1$, where the PS behaves as $k^{-4}$ with superimposed oscillations [cf. Eq. (3.28)]. These wave numbers correspond to $k_{0}\lesssim k\lesssim k_{\mathrm{eq}}$ which re-enter the Hubble radius during the epoch of matter domination.
@@ 4 Regularization of the PS of PGWs
In the previous section, assuming the Bunch-Davies initial conditions, we calculated the PS of PGWs at the time of radiation-matter equality as well as today. We found that, over small scales such that $y_{\mathrm{e}}\gg 1$, the PS behaves as $k^{2}$. As is well known, it is such a rise that leads to ultraviolet divergences in the coincident limit of the two-point functions of quantum fields in real space. In quantum field theory, these divergences are usually dealt with the help of regularization and renormalization procedures (for discussions in this context, see the books [104, 105, 106, 107]). For this reason, we shall also regularize the two-point function in Fourier space, i.e. the PS (for related discussions, see Refs. [84, 91, 85, 86, 87, 88, 89, 90]). We should emphasize that this last statement is non-trivial. Note that, in real space, divergences arise in the coincident limit of two-point functions of quantum fields because they involve an integral over all the wave numbers. In contrast, the two-point function in Fourier space, i.e. the PS, is evaluated at a given wave number. Therefore, actually, the PS does not contain any divergence. Nevertheless, the PS has to be regularized in order to ensure that the unphysical $k^{2}$ behavior does not arise at small scales.
@@ 4.1 Employing the method of adiabatic subtraction
The regularization procedure that is ideally suited for time-dependent situations such as the FLRW universe is the method of adiabatic regularization (for the original discussions, see Refs. [108, 109, 110, 111, 112, 113]; in this context, also see the books [106, 107]). Since we shall be interested in regularizing quantities at a given wave number, the procedure turns out to be relatively simple to employ [^4] In due course, it will be interesting to examine if other regularization procedures, such as, say, the point-splitting method, lead to the same results.. The method involves identifying the contribution to the two-point function at a given adiabatic order and subtracting it from the complete contribution. In order to parametrize the slowness of the changes in the background and calculate the corresponding contributions, it proves to be convenient to introduce a parameter, say, $T$, often referred to as the adiabatic parameter. Thereafter, an expansion in the inverse powers of $T$ is carried out and the terms at the order $T^{-n}$ is referred to as the contributions at $n$th adiabatic order [106]. Eventually, the parameter $T$ is set to unity at the end of the calculations. As we shall see, the regularization of the PS is carried out by subtracting terms up to the second adiabatic order.
@@
where $\omega_{k}^{2}(\eta)$ and $\sigma^{2}(\eta)$ are functions of time. For instance, in the case of a massive scalar field, the quantity $\omega_{k}^{2}(\eta)$ will depend on the wave number, the scale factor of the FLRW universe and its time derivatives, while the quantity $\sigma^{2}(\eta)$ will depend on the mass of the field and the scale factor. To solve such an equation, let us assume the WKB ansatz
@@
According to Eq. (4.4), we have (when $T$ eventually set to unity)
@@ 4.2 Regularization of the PS of PGWs
To regularize the PS of PGWs, we need to calculate the quantity $|\mu_{k}|_{\mathrm{ad}}^{2}$ up to the second adiabatic order. We have
@@ 4.3 Regularized PS in the case of the instantaneous transitions
Figure 4: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the radiation-dominated epoch at $\eta=-2\eta_{\mathrm{e}}=2|\eta_{\mathrm{e}}|$ have been plotted (in red and blue) for the case of instantaneous transition from de Sitter inflation [cf. Eq. (4.17)]. We have plotted the PS for the same value of $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}$ as in the previous figure. Note that, over $y_{\mathrm{e}}>1$ (indicated by the vertical dotted line), while the original PS rises as $k^{2}$ (with oscillations superimposed upon it), the regularized PS oscillates about zero with a constant amplitude. In other words, over $k\gtrsim k_{\mathrm{e}}$, $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ can turn negative.
@@
Figure 5: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated today at the conformal time $\eta_{0}$, have been plotted (in red and blue) for the case of instantaneous transition from de Sitter inflation to the epochs of radiation and matter domination [cf. Eqs. (3.2) and (4.2)]. We have chosen to work with $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}=1.8\times 10^{16}\,\mathrm{GeV}$ as in the previous two figures. Also, as in Fig. 3, the vertical lines indicate the wave numbers $k_{0}/a_{0}$, $k_{\mathrm{eq}}/a_{0}$ and $k_{\mathrm{e}}/a_{0}$. Moreover, as in Fig. 3, we have included the sensitivity curves of the various GW observatories. It is clear that the process of regularization truncates the $k^{2}$ rise for $k\gtrsim k_{\mathrm{e}}$. Also, note that the regularization procedure does not alter the PS over scales $k\lesssim k_{\mathrm{e}}$. We find that, over $k\gtrsim k_{\mathrm{e}}$, the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ oscillates about zero as illustrated in the previous figure. However, the amplitude of the oscillations is significantly suppressed during the late stages of matter domination when compared to the values during the early stages of radiation domination.
@@ 4.4 Adiabatic regularization in real space
The fact that the regularized PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ is negative over some scales (as we saw in the previous section) may be viewed as problematic and, in this section, we shall discuss this issue in some detail. To begin with, note that the original PS $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ is a positive definite quantity, as should be evident from Eq. (2.11). However, since the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ is the difference between two quantities, it does not have to be positive definite. Hence, in principle, there is no reason to expect $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ to remain positive on all scales. Secondly, it is the two-point correlation function of PGWs in real space [cf. Eq. (2.10)] that is an observable and, therefore, physical requirements should be imposed only on this quantity. The PS can be compared to the sensitivities of various GW detectors, as we have done, for instance, in Fig. 3. But, the PS of PGWs is only a derived quantity and it is inferred from the physical constraints arrived at on the two-point correlation function in real space. These constraints can imply a non-standard behavior for the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$. However, provided the corresponding correlation function in real space behaves in a proper fashion, the non-standard behavior of $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ need not be viewed as problematic.
Let us now highlight the reasons as to why it is the two-point correlation function in real space is the quantity that can be observationally probed. At a given spatial location, say, ${\bm{x}}=0$, the signal, say, $\mathcal{S}(t)$, observed by a GW detector can be expressed as [97, 98]
@@
As we have discussed before, to regularize the PS, it is the contribution at the second adiabatic order that is to be subtracted from the complete PS. In real space, the contribution to the two-point correlation function of the tensor perturbations due to the second adiabatic order in Fourier space [cf. Eq. (4.11)] can be expressed as
@@
We can evaluate the second integral by introducing a comoving infrared cutoff, say, $k_{{}_{\mathrm{IR}}}$. Upon doing so, we finally arrive at the adiabatic contribution at the second order to the two-point function in real space to be
@@
Finally, a set of last remarks are in order. Often, it has been suggested that the signal of GWs has to be averaged over time (in this context, see, for instance, Ref. [98]). We find that, if we average over time, the PS of PGWs, evaluated during the epoch of radiation domination [cf. Eq. (3.5)], reduces to the following simple form:
@@ 5 Smoothing the transition from inflation to radiation domination
$U(\eta)=\left\{\begin{array}[]{ll}\!\frac{q\,(q-1)}{\eta^{2}},&\mbox{for}\penalty 10000\ \eta\leq\eta_{\mathrm{e}},\\ \!\frac{2}{\eta_{\mathrm{e}}^{2}\,\Delta\eta}\!\left\{\left[\frac{\gamma(w)}{(1-\Delta\eta/|\eta_{\mathrm{e}}|+\eta_{w}/|\eta_{\mathrm{e}}|)^{2}}-\frac{q\,(q-1)}{2}\right](\eta-\eta_{\mathrm{e}})+\frac{q\,(q-1)}{2}\,\Delta\eta\right\},&\mbox{for}\penalty 10000\ \eta_{\mathrm{e}}\leq\eta\leq\eta_{\mathrm{e}}+\Delta\eta,\\ \!\frac{2\,\gamma(w)}{(\eta-\eta_{w})^{2}},&\mbox{for}\penalty 10000\ \eta\geq\eta_{\mathrm{e}}+\Delta\eta.\end{array}\right$. (5.4)
@@
Figure 6: The behavior of the quantity $U(\eta)=a^{\prime\prime}/a$ has been plotted in the case of transition from de Sitter inflation (corresponding to q=1q=-1) to the epoch of radiation domination (corresponding to $w=1/3$), when $U(\eta)$ has been smoothed with a linear function [cf. Eq. (5.4)]. Note that we have set $\Delta\eta=2|\eta_{\mathrm{e}}|$ in plotting the figure. We have also indicated the beginning and the end of the smooth transition (as vertical dotted and dashed lines, respectively).
@@ 6 Evolution of PGWs across the smoother transition
Having obtained the solution for the background across the smooth linear transition, let us now construct the solution for the tensor mode functions.
@@ 6.1 Evolution during the smoother transition
We are now in a position to solve the equation (2.9) for the rescaled mode function $\mu_{k}(\eta)$. The solutions during inflation and the radiation dominated era are already known and, hence, let us focus on the solution during the transition. If we define
@@
which can be recognized again as the Airy differential equation. As a result, the general solution for the rescaled mode function can be expressed as
@@
where $T^{\prime}_{1/2-q}(y)$ denotes the derivative of the function $T_{1/2-q}(y)$ with respect to its argument. Note that the mode functions $i_{k}(\eta)$ and $j_{k}(\eta)$ are described by the Airy functions with the argument $\tau(\eta)$ [see Eq. (6.3)]. The quantity $\tau_{\mathrm{e}}$ can be expressed as
@@
and, in the limit $y_{\mathrm{e}}\gg 1$, $\tau(\eta_{\mathrm{e}})$ is a large, negative quantity. As a consequence, it is convenient to use the following exact expressions for the Airy functions [114]:
@@
where $\varepsilon=\pm 1$ depending on the value of the equation-of-state parameter $w$ in the subsequent epoch [cf. Eq. (5.16)]. (Recall that $\varepsilon=-1$ in the case of transition to a radiation dominated epoch.) Finally, on using the expressions (6.5) for the Wronskian and the inflationary mode function (6.6), we find that the coefficient $B_{1}$ can be written as

### v2

@@ 3.2 PS during matter domination
Since $a_{0}$ and $H_{0}$ are the scale factor and the Hubble parameter today, $k_{0}/a_{0}=H_{0}$ is the smallest wave number of observational interest. On substituting the above expressions for $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ in the PS (3.16) and considering the limit $z\gg 1$, which corresponds to $k_{0}\ll k\ll k_{\mathrm{eq}}$, we obtain that
@@
Figure 2: The PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ evaluated today, i.e. at $\eta_{0}$, has been plotted for the case of instantaneous transitions from de Sitter inflation to the epoch of radiation domination and subsequently to the epoch of matter domination [cf. Eq. (3.2)]. We have plotted the PS for the three values of $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}$ we had plotted in the previous figure. Note that, in addition to the different behavior of the PS in the three regimes mentioned in the previous figure, there arises a fourth regime when $x_{\mathrm{eq}}\ll 1$ and $z\gg 1$, where the PS behaves as $k^{-4}$ with superimposed oscillations [cf. Eq. (3.28)]. These wave numbers correspond to $k_{0}\lesssim k\lesssim k_{\mathrm{eq}}$ which re-enter the Hubble radius during the epoch of matter domination.
@@ 4 Regularization of the PS of PGWs
In the previous section, assuming the Bunch-Davies initial conditions, we calculated the PS of PGWs at the time of radiation-matter equality as well as today. We found that, over small scales such that $y_{\mathrm{e}}\gg 1$, the PS behaves as $k^{2}$. As is well known, it is such a rise that leads to ultraviolet divergences in the coincident limit of the two-point functions of quantum fields in real space. In quantum field theory, these divergences are usually dealt with the help of regularization and renormalization procedures (for discussions in this context, see the books [132, 134, 40, 110]). For this reason, we shall also regularize the two-point function in Fourier space, i.e. the PS (for related discussions, see Refs. [112, 69, 20, 21, 22, 128, 58, 114]). We should emphasize that this last statement is non-trivial. Note that, in real space, divergences arise in the coincident limit of two-point functions of quantum fields because they involve an integral over all the wave numbers. In contrast, the two-point function of quantum fields in Fourier space, i.e. the PS, is evaluated at a given wave number. Therefore, actually, the PS does not contain any divergence. Nevertheless, the PS has to be regularized in order to ensure that the unphysical $k^{2}$ behavior does not arise at small scales.
@@ 4.1 Employing the method of adiabatic subtraction
There exist a variety of regularization procedures to handle the divergences that are encountered when examining the behavior of quantum fields in curved spacetimes (in this context, see the books [40, 110]). These include dimensional regularization (see, for instance, Refs. [43, 134]), zeta-function regularization (for the initial discussions, see Refs. [65, 64, 83]; for specific applications, see Ref. [66]), proper-time regularization (for the original discussion, see Ref. [122]; for additional discussions in electromagnetic backgrounds, see Refs. [61, 60]; for early applications in curved spacetimes, see, for instance, Refs. [59, 65]; for further discussions, see Refs. [40, 110]), point-splitting regularization (for the initial discussions, see Refs. [50, 56, 55, 44, 129, 130, 46, 45, 51]; for further discussions, see Refs. [40, 110]), Hadamard regularization (see, for example, Refs. [126, 57] and, for a discussion related to the topic of our interest, see Refs. [106, 107]), and adiabatic regularization (for the original discussions, see Refs. [72, 111, 73, 47, 48, 25]). Among these regularization procedures, the approach that is the most convenient to implement in time-dependent situations such as the FLRW universe is the method of adiabatic regularization. Since we shall be interested in regularizing quantities at a given wave number, the procedure turns out to be relatively simple to employ. We should also mention that the different methods of regularization are expected to lead to the same results. In due course, it will be interesting to reproduce the results we shall obtain below using, say, the method of point-splitting.
The method of adiabatic regularization involves identifying the contribution to the two-point function at a given adiabatic order and subtracting it from the complete contribution. In order to parametrize the slowness of the changes in the background and calculate the corresponding contributions, it proves to be convenient to introduce a parameter, say, $T$, often referred to as the adiabatic parameter. Thereafter, an expansion in the inverse powers of $T$ is carried out and the terms at the order $T^{-n}$ is referred to as the contributions at $n$th adiabatic order [40]. Eventually, the parameter $T$ is set to unity at the end of the calculations. As we shall see, the regularization of the PS is carried out by subtracting terms up to the second adiabatic order.
@@
where $\omega_{k}^{2}(\eta)$ and $\sigma^{2}(\eta)$ are functions of time. For instance, in the case of a massive scalar field, the quantity $\omega_{k}^{2}(\eta)$ will depend on the wave number, the mass of the field and the scale factor of the FLRW universe, while the quantity $\sigma^{2}(\eta)$ will depend on the scale factor and its time derivatives. To solve such an equation, let us assume the WKB ansatz
@@
According to Eq. (4.4), when $T$ is eventually set to unity, we have
@@ 4.2 Regularization of the PS of PGWs
To regularize the PS of PGWs, we need to calculate the quantity $|\mu_{k}^{\mathrm{ad}}|^{2}$ up to the second adiabatic order. We have
@@ 4.3 Regularized PS in the case of the instantaneous transitions
Figure 4: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the radiation-dominated epoch at $\eta=-2\eta_{\mathrm{e}}=2|\eta_{\mathrm{e}}|$ have been plotted (in red and blue) for the case of instantaneous transition from de Sitter inflation [cf. Eq. (4.17)]. We have plotted the PS for the same value of $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}$ as in the previous figure. Note that, over $y_{\mathrm{e}}>1$ (demarcated by the vertical dotted line), while the original PS rises as $k^{2}$ (with oscillations superimposed upon it), the regularized PS oscillates about zero with a constant amplitude. In other words, over $k\gtrsim k_{\mathrm{e}}$, $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ can turn negative.
@@
Figure 5: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated today at the conformal time $\eta_{0}$, have been plotted (in red and blue) as a function of frequency $f$ for the case of instantaneous transitions from de Sitter inflation to the epochs of radiation and matter domination [cf. Eqs. (3.2) and (4.2)]. We have chosen to work with $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}=1.8\times 10^{16}\,\mathrm{GeV}$ as in the previous two figures. Also, as in Fig. 3, the vertical lines indicate the frequencies corresponding to the wave numbers $k_{0}/a_{0}$, $k_{\mathrm{eq}}/a_{0}$ and $k_{\mathrm{e}}/a_{0}$. Moreover, as in Fig. 3, we have included the sensitivity curves of the various GW observatories. It is clear that the process of regularization truncates the $k^{2}$ rise for $k\gtrsim k_{\mathrm{e}}$. Also, note that the regularization procedure does not alter the PS over scales $k\lesssim k_{\mathrm{e}}$. We find that, over $k\gtrsim k_{\mathrm{e}}$, the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ oscillates about zero as illustrated in the previous figure. However, the amplitude of the oscillations is significantly suppressed during the late stages of matter domination when compared to the values during the early stages of radiation domination.
@@ 4.4 Adiabatic regularization in real space
The fact that the regularized PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ is negative over some scales (as we saw in the previous section) may be viewed as problematic and, in this section, we shall discuss this issue in some detail. To begin with, note that the original PS $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ is a positive definite quantity, as should be evident from Eq. (2.11). However, since the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ is the difference between two quantities, it does not have to be positive definite. Hence, in principle, there is no reason to expect that $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ will remain positive on all scales. Secondly, it is the two-point correlation function of PGWs in real space [cf. Eq. (2.10)] that is an observable and, therefore, physical requirements should be imposed only on this quantity. The PS can be compared to the sensitivities of various GW detectors, as we have done, for instance, in Fig. 3. But, the PS of PGWs is only a derived quantity and it is inferred from the physical constraints arrived at on the two-point correlation function in real space. These constraints can imply a non-standard behavior for the regularized PS $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$. However, provided the corresponding correlation function in real space behaves in a proper fashion, the non-standard behavior of $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$ need not be viewed as problematic.
Let us now highlight the reasons as to why the two-point correlation function in real space is the quantity that can be observationally probed. At a given spatial location, say, ${\bm{x}}=0$, the signal, say, $\mathcal{S}(t)$, observed by a GW detector can be expressed as [96, 97]
@@
As we have discussed before, to regularize the PS, it is the contribution up to the second adiabatic order that is to be subtracted from the complete PS. In real space, the contribution to the two-point correlation function of the tensor perturbations up to the second adiabatic order in Fourier space [cf. Eq. (4.11)] can be expressed as
@@
We can evaluate the second integral by introducing a comoving infrared cutoff, say, $k_{{}_{\mathrm{IR}}}$. Upon doing so, we finally arrive at the adiabatic contribution up to the second order to the two-point function in real space to be
@@
Finally, a last set of remarks are in order. Often, it has been suggested that the signal of GWs has to be averaged over time (in this context, see, for instance, Ref. [97]). We find that, if we average over time, the PS of PGWs, evaluated during the epoch of radiation domination [cf. Eq. (3.5)], reduces to the following simple form:
@@ 5 Smoothing the transition from inflation to radiation domination
$U(\eta)=\left\{\begin{array}[]{ll}\!\frac{q\,(q-1)}{\eta^{2}},&\mbox{for}\penalty 10000\ \eta\leq\eta_{\mathrm{e}},\\ \!\frac{2}{\eta_{\mathrm{e}}^{2}\,\Delta\eta}\!\left\{\left[\frac{\gamma(w)}{(1-\Delta\eta/|\eta_{\mathrm{e}}|+\eta_{w}/|\eta_{\mathrm{e}}|)^{2}}-\frac{q\,(q-1)}{2}\right](\eta-\eta_{\mathrm{e}})+\frac{q\,(q-1)}{2}\,\Delta\eta\right\},&\mbox{for}\penalty 10000\ \eta_{\mathrm{e}}\leq\eta\leq\eta_{\mathrm{e}}+\Delta\eta,\\ \!\frac{2\,\gamma(w)}{(\eta-\eta_{w})^{2}},&\mbox{for}\penalty 10000\ \eta\geq\eta_{\mathrm{e}}+\Delta\eta,\end{array}\right$. (5.4)
where $\gamma(w)$ is given by
$\gamma(w)=\frac{(1-3w)}{(1+3w)^{2}}.$ (5.5)
@@
Figure 6: The behavior of the ‘effective potential’ $U(\eta)=a^{\prime\prime}/a$ has been plotted in the case of transition from de Sitter inflation (corresponding to q=1q=-1) to the epoch of radiation domination (corresponding to $w=1/3$), when $U(\eta)$ has been smoothed with a linear function [cf. Eq. (5.4)]. Note that we have set $\Delta\eta=2|\eta_{\mathrm{e}}|$ in plotting the figure. We have also indicated the beginning and the end of the smooth transition (as vertical dotted and dashed lines, respectively).
@@ 6 Evolution of PGWs across the smoother transition
Having obtained the solution for the background across the smoother linear transition, let us now construct the solution for the tensor mode functions.
@@ 6.1 Evolution during the smoother transition
We are now in a position to solve Eq. (2.9) for the rescaled mode function $\mu_{k}(\eta)$. The solutions during inflation and the radiation-dominated era are already known and, hence, let us focus on the solution during the transition. If we define
@@
which can be recognized again as the Airy differential equation. As a result, the general solution for the rescaled mode function can be expressed as [78]
@@
where $T^{\prime}_{1/2-q}(y)$ denotes the derivative of the function $T_{1/2-q}(y)$ with respect to its argument. Note that the mode functions $i_{k}(\eta)$ and $j_{k}(\eta)$ are described by the Airy functions with the argument $\tau(\eta)$ [see Eq. (6.3)]. The quantity $\tau(\eta_{\mathrm{e}})\equiv\tau_{\mathrm{e}}$ can be expressed as
@@
and, in the limit $y_{\mathrm{e}}\gg 1$, $\tau_{\mathrm{e}}$ is a large, negative quantity. As a consequence, it is convenient to use the following exact expressions for the Airy functions [12]:
@@
where $\varepsilon=\pm 1$ depending on the value of the equation-of-state parameter $w$ in the subsequent epoch [cf. Eq. (5.17)]. (Recall that $\varepsilon=-1$ in the case of transition to a radiation-dominated epoch.) Finally, on using the expressions (6.5) for the Wronskian and the inflationary mode function (6.6), we find that the coefficient $B_{1}$ can be written as